# BARK: A Fully Bayesian Tree Kernel for Black-box Optimization

### v1

@@ Abstract
Machine Learning, ICML
@@ 2 Related Work
Bayesian optimization over mixed feature spaces has been extensively researched in recent years. Techniques include using RKHS embeddings (Buathong et al., 2020), diffusion kernels (Deshwal et al., 2021), multi-armed bandit style optimization (Ru et al., 2020), and a plethora of other methods (Garrido-Merchán & Hernández-Lobato, 2020; Häse et al., 2021; Daxberger et al., 2021; Daulton et al., 2022; Deshwal et al., 2023) and benchmarks (Dreczkowski et al., 2023). This surge in interest also includes the provision of software packages for mixed feature BO such as SMAC (Lindauer et al., 2022) and scikit-optimize (Head et al., 2018). More specifically, there are also several approaches using tree-based models for black-box optimization (Ammari et al., 2023; Mistry et al., 2021; Thebelt et al., 2021, 2022b). Jenatton et al. (2017) exploit a known tree structure for more efficient BO. Lei et al. (2021) use BART for BO, however, they do not fully exploit the implicit GP structure in the model as we do. Finally, Thebelt et al. (2022a) show how to use GPs with tree kernels for BO with optimal acquisition function optimization, however, they fit the trees using a two-step procedure.
@@ 5.2 Differences from the BART model
Crucially, our kernel perspective means that we never sample leaf values - the marginal likelihood integrates over all the values of the leaves. In this way, BARK samples distributions, which can then be used to build acquisition functions, motivating our use of the GP likelihood. In contrast, BART samples the leaf values sequentially, hence uses the likelihood conditional on leaf values. BART is far less sample efficient when approximating the posterior leaf distribution. Both BART and BARK have the same probabilistic model: marginalizing the BART likelihood over leaf values yields the BARK likelihood, as in Eq. 4.
BARK proposals. In BART, splitting points are typically uniformly sampled from the data seen by a given node (Chipman et al., 2010; Kapelner & Bleich, 2016). This choice is made such that any leaf node will contain a nonempty subset of the domain (no ‘logically empty’ leaves). However, this approach performs poorly in the BO setting, since finding potential future queries requires good uncertainty estimation far from data (see Section 6). Instead, we sample the splitting rule uniformly from the domain. To avoid creating logically empty nodes, each splitting rule is sampled from the subspace of the domain that reaches that node. For example, if $x\in[0,1]$, and the root splitting rule is $x>0.5$, then the splitting value of the left child is be sampled from $[0,0.5]$. For categorical features with $S$ categories that reach node a node, the splitting rule is sampled uniformly from $P(S)\setminus\{S,\{\}\}$, i.e., the power set of $S$ excluding trivial splits. This enables greater sample efficiency than a one-hot encoding of categories.
BARK cannot sample directly from the posterior over $\sigma_{y}^{2}$, and so must sample the noise using MCMC. To enforce non-negativity, a proposal is generated by taking a Gaussian walk in an unconstrained space, using the softplus transform. Details for the BARK proposals are given in Appendix E.
@@ 6.1 Optimizing the acquisition function
Typical BO approaches use gradient-based methods to optimize the acquisition function (AF), such as L-BFGS (Zhu et al., 1997). However, our piecewise constant prior has zero gradient almost everywhere. We turn to mixed integer linear programming to define the optimization problem. We use the framework provided by Thebelt et al. (2022a); Mišić (2020) to encode the tree structure using constraints on binary variables. The full formulation is given in Appendix G. This has the additional benefit of global optimization, where gradient-based methods can be stuck in local optima.
@@ 6.3 The Laplace approximation for regret bounds
The functions in the BARK prior are discontinuous and non-stationary. This violates the typical assumptions required for regret bounds in the Bayesian optimization literature (Srinivas et al., 2012). Previous work shows that, under some strong assumptions, the tree kernel can be approximated as a Matern-$\frac{1}{2}$ kernel (Linero, 2017), which in turn admits sub-linear regret bounds of $\tilde{O}(T^{\frac{1+2D}{2+2D}})$, where $D$ is the dimension of the domain (Wang et al., 2023b). Appendix F provides further discussion on the quality of this approximation.
@@ 7 Experiments
Baselines. We compare BARK to several baselines. For each baseline, we use the upper confidence bound (UCB) AF. BART uses the Chipman et al. (2010) surrogate, evaluated on a grid, and uses function samples to estimate the UCB (Wilson et al., 2018). GP-RBF uses the squared-exponential kernel, with the default priors from Hvarfner et al. (2024). For enumerable mixed feature spaces, we use a linear combination of sum and product kernels, as in Ru et al. (2020). For domains where the number of combinations of categories makes enumeration infeasible, we use a continuous relaxation, first applying a one-hot encoding to categorical features. LeafGP uses a two-step forest kernel, where the forest is fit using gradient boosted trees, then used as a kernel in a GP (Thebelt et al., 2022a). This model can be used in a global optimization of the UCB acquisition function. SMAC uses an ensemble of trees, using sample variance to measure uncertainty (Lindauer et al., 2022). Entmoot defines a mean function with a tree surrogate, and uses a distance-based uncertainty (Thebelt et al., 2021). Finally, BARKPrior samples tree kernels from the prior, without performing the MCMC steps to sample the posterior. Appendix B provides further details of the methods.
@@ 7.1 Model fit comparison
We perform regression with both BART and BARK. We use the default setting for the model hyperparameters; for discussion on the selection of these values, see Appendix D.
@@
We use real datasets from the UCI Repository (Cortez & Silva, 2008; Nash et al., 1994; Quinlan, 1993; Yeh, 1998; Dua & Graff, 2017). We compare the predictive performances in Section 7.1. The performance of the two models is similar, suggesting that BARK can achieve similarly strong regression performance to BART. Our slight increase in NLPD performance may be due to our GP perspective, which effectively allows us to marginalize over the distribution of leaf values. Appendix C discusses the training time on these regression problems.
@@ 7.2 Synthetic Benchmarks
TreeFunction and TreeFunctionCat are functions sampled from the BART prior, each with 10 continuous dimensions, and the latter with an additional 10 categorical dimensions. DiscreteAckley and DiscreteRosenbrock are partially discretized functions from Dreczkowski et al. (2023); Bliek et al. (2021). Appendix B has further benchmark details.
Figure 5: Optimization of synthetic benchmarks.
@@
DiscreteRosenbrock shows a failure case of BARK; the optimum lies in a very shallow region, which the piecewise-constant prior of BARK models worse than Entmoot’s distanced-based uncertainty quantification.
@@ 7.3 Applied benchmarks
GP-RBF achieves the best final value on XGBoostMNIST. When it is feasible to enumerate over every categorical combination, GP-RBF is a strong choice. However, when continuous relaxation is used due to the highly combinatorial nature of the domain, GP-RBF is more likely to be stuck in local optima, and perform worse than BARK.
Figure 6: Optimization of applied benchmarks. XGBoostMNIST allows enumeration of GP-RBF, whereas other benchmarks require continuous relaxation.
@@ Acknowledgments
Funding for this work was provided by BASF SE, Ludwigshafen am Rhein, and EPSRC through the Modern Statistics and Statistical Machine Learning CDT for TB (EP/Y034813/1) and JPF (EP/S023151/1). TB is also funded by the EPSRC IConIC Prosperity Partnership (EP/X025292/1).
Appendix A Symbols
@@
Appendix B Experimental details
B.1 Method details
BARK. For both model fitting and BO, we use 16 samples for the BARK kernel. Increasing the number of samples would (marginally) improve the model fit, however this would increase the complexity of the optimization problem, meaning that the optimization would be less likely to converge to the optimum of the acquisition function. BARK uses 1000 burn-in samples and 400 kernel samples, running in parallel with 4 chains. We use a thinning rate of 100 to obtain the 16 samples. We evaluate the effect of increasing the number of samples in Appendix D.
@@ B.1 Method details
There is no procedure in the literature for optimizing acquisition functions for BART, and the MIP formulation would be too expensive for the 1000s of BART samples required. We therefore evaluate the surrogate on a grid of $\max(2^{5D},2^{14})$ points using a space-filling approach (Sobol, 1967). At each iteration, we sample a new grid to better explore the domain.
@@
We use the BoFire v0.0.16 implementation (Dürholt et al., 2024), which in turn uses BoTorch v0.11.3 (Balandat et al., 2020).
@@
B.2 Benchmark details
A description of the benchmarks used in Bayesian optimization is given in Tables 3 and 4. For each BO benchmark, we generate a set of $\min(2D,30)$ initial points, and run the optimization loop for 100 iterations, reporting the minimum value observed up to iteration $t$.
@@ B.2 Benchmark details
Appendix C Sampling time for BARK and BART
BART and BARK both take significantly longer to train than the other baselines, as they are MCMC-based methods. Given $N$ observations, training BART has $O(N)$ complexity, whereas BARK trains with $O(N^{3})$ complexity (which is reduced to $O(N^{2})$ when sampling tree structures due to the low-rank update discussed in Section 5). We therefore expect BART to scale far more favorably with many observations. In Appendix C, we show the time taken to train the models on the regression problems from Section 7. The fitting time for BARK is reasonable in the low-data regime of BO, matching that of BART when only 40 training points are used. Since we apply our model in the BO setting, where the evaluation of the black-box function may take hours, or possibly even days, we argue that this time taken is not a significant limitation of the model. We also note that we use fewer samples for BARK (400 samples per parallel chain) compared to BART (1000 samples per chain).

### v2

@@ Abstract
Bayesian optimization, tree kernel
@@ 1 Introduction
We motivate using tree-kernels by exploiting the strong modeling power of tree-based approaches, which can be used for mixed input spaces. Moreover, we can formulate acquisition functions with these tree models as optimization problems that can be solved globally, where gradient-based methods often struggle with discrete inputs.
@@ 2 Related Work
Bayesian optimization over mixed feature spaces has been extensively researched in recent years. Techniques include using RKHS embeddings (Buathong et al., 2020), diffusion kernels (Deshwal et al., 2021), multi-armed bandit style optimization (Ru et al., 2020), and a plethora of other methods (Garrido-Merchán & Hernández-Lobato, 2020; Häse et al., 2021; Daxberger et al., 2021; Daulton et al., 2022; Deshwal et al., 2023; Wan et al., 2021) and benchmarks (Dreczkowski et al., 2023). This surge in interest also includes the provision of software packages for mixed feature BO such as SMAC (Lindauer et al., 2022) and scikit-optimize (Head et al., 2018). More specifically, there are also several approaches using tree-based models for black-box optimization (Ammari et al., 2023; Mistry et al., 2021; Thebelt et al., 2021, 2022b). Jenatton et al. (2017) exploit a known tree structure for more efficient BO. Lei et al. (2021) use BART for BO, however, they do not fully exploit the implicit GP structure in the model as we do. Finally, Thebelt et al. (2022a) show how to use GPs with tree kernels for BO with optimal acquisition function optimization, however, they fit the trees using a two-step procedure.
@@ 5.2 Differences from the BART model
Crucially, our kernel perspective means that we never sample leaf values - the marginal likelihood integrates over all the values of the leaves. In this way, BARK samples distributions, which can then be used to build acquisition functions, motivating our use of the GP likelihood. In contrast, BART samples the leaf values sequentially, hence uses the likelihood conditional on leaf values. In order to approximate the posterior leaf distribution, BART requires a large number of function samples, which in turn means that it is infeasible to formulate an optimization problem over the BART acquisition function. Both BART and BARK have the same probabilistic model: marginalizing the BART likelihood over leaf values yields the BARK likelihood, as in Eq. 4.
BARK proposals. In BART, splitting points are typically uniformly sampled from the data seen by a given node (Chipman et al., 2010; Kapelner & Bleich, 2016). This choice is made such that any leaf node will contain a nonempty subset of the domain (no ‘logically empty’ leaves). However, this approach performs poorly in the BO setting, since finding potential future queries requires good uncertainty estimation far from data (see Section 6). Instead, we sample the splitting rule uniformly from the domain. To avoid creating logically empty nodes, each splitting rule is sampled from the subspace of the domain that reaches that node. For example, if $x\in[0,1]$, and the root splitting rule is $x>0.5$, then the splitting value of the left child is be sampled from $[0,0.5]$. For categorical features with $S$ categories that reach node a node, the splitting rule is sampled uniformly from $P(S)\setminus\{S,\{\}\}$, i.e., the power set of $S$ excluding trivial splits. This enables greater sample efficiency than a one-hot encoding of categories. We investigate the impact of this modeling choice in Appendix B for toy regression examples.
BARK cannot sample directly from the posterior over $\sigma_{y}^{2}$, and so must sample the noise using MCMC. To enforce non-negativity, a proposal is generated by taking a Gaussian walk in an unconstrained space, using the softplus transform. Details for the BARK proposals are given in Appendix H.
@@ 6.1 Optimizing the acquisition function
Typical BO approaches use gradient-based methods to optimize the acquisition function (AF), such as L-BFGS (Zhu et al., 1997). However, our piecewise constant prior has zero gradient almost everywhere. We turn to mixed integer linear programming to define the optimization problem. We use the framework provided by Thebelt et al. (2022a); Mišić (2020) to encode the tree structure using constraints on binary variables. The full formulation is given in Appendix J. This has the additional benefit of global optimization, where gradient-based methods can be stuck in local optima.
@@ 6.3 The Laplace approximation for regret bounds
The functions in the BARK prior are discontinuous and non-stationary. This violates the typical assumptions required for regret bounds in the Bayesian optimization literature (Srinivas et al., 2012). Previous work shows that, under some strong assumptions, the tree kernel can be approximated as a Matern-$\frac{1}{2}$ kernel (Linero, 2017), which in turn admits sub-linear regret bounds of $\tilde{O}(T^{\frac{1+2D}{2+2D}})$, where $D$ is the dimension of the domain (Wang et al., 2023b). Appendix I provides further discussion on the quality of this approximation.
@@ 7 Experiments
Baselines. We compare BARK to several baselines. For each baseline, we use the upper confidence bound (UCB) AF. BART uses the Chipman et al. (2010) surrogate, evaluated on a grid, and uses function samples to estimate the UCB (Wilson et al., 2018). GP-RBF uses the squared-exponential kernel, with the default priors from Hvarfner et al. (2024). For enumerable mixed feature spaces, we use a linear combination of sum and product kernels, as in Ru et al. (2020). To optimize over mixed feature spaces, we use an alternating approach to perform optimization, similar to (Wan et al., 2021), where we alternate between optimizing the continuous parameters with gradient-based methods, and exploring local perturbations in the discrete parameters. For domains where the number of combinations of categories is low, we also include an enumeration approach, where each unique combination of discrete values is fixed and many parallel continuous optimizations are performed. LeafGP uses a two-step forest kernel, where the forest is fit using gradient boosted trees, then used as a kernel in a GP (Thebelt et al., 2022a). This model can be used in a global optimization of the UCB acquisition function. SMAC uses an ensemble of trees, using sample variance to measure uncertainty (Lindauer et al., 2022). Entmoot defines a mean function with a tree surrogate, and uses a distance-based uncertainty (Thebelt et al., 2021). Finally, BARKPrior samples tree kernels from the prior, without performing the MCMC steps to sample the posterior. Appendix D provides further details of the methods.
@@ 7.1 Model fit comparison
We perform regression with both BART and BARK. The purpose of this section is to demonstrate that we do not lose predictive power when using the kernel perspective of the BARK model. We show this in the tabular regression setting, where BART is already known to be strong. We use the default setting for the model hyperparameters; for discussion on the selection of these values, see Appendix G.
@@
We use real datasets from the UCI Repository (Cortez & Silva, 2008; Nash et al., 1994; Quinlan, 1993; Yeh, 1998; Dua & Graff, 2017). We compare the predictive performances in Section 7.1. The performance of the two models is similar, suggesting that BARK can achieve similarly strong regression performance to BART. Our slight increase in NLPD performance may be due to our GP perspective, which effectively allows us to marginalize over the distribution of leaf values. Appendix F discusses the training time on these regression problems, and Section E.1 provides a comparison to the GP-RBF model.
@@ 7.2 Synthetic benchmarks
TreeFunction and TreeFunctionCat are functions sampled from the BART prior, each with 10 continuous dimensions, and the latter with an additional 10 categorical dimensions. DiscreteAckley and DiscreteRosenbrock are partially discretized functions from Dreczkowski et al. (2023); Bliek et al. (2021). Appendix D has further benchmark details.
Figure 5: Optimization of synthetic benchmarks. The shaded regions contain the 25th and 75th percentile of regret achieved across the 20 runs.
@@
DiscreteRosenbrock shows a failure case of BARK; the optimum lies in a very shallow region, which the piecewise-constant prior of BARK models worse than the smooth prior of the GP-RBF, and Entmoot’s distanced-based uncertainty quantification. These synthetic benchmarks tend to favor the smoothness assumptions of the RBF kernel over the stepwise-constant functions in the BARK posterior. We provide further results for continuous synthetic functions in Section E.2.
@@ 7.3 Applied benchmarks
GP-RBF achieves the best final value on XGBoostMNIST. When it is feasible to enumerate over every categorical combination, GP-RBF is a strong choice. However, in the other applied benchmarks where the alternating optimization approach must be employed, BARK is highly competitive with GP-RBF.
Figure 6: Optimization of applied benchmarks. XGBoostMNIST allows enumeration of GP-RBF, whereas other benchmarks require alternating optimization.
@@
7.4 Wall-time discussion
BARK’s most significant limitation is its wall-time performance. Fitting the model takes approximately 50 seconds, and the optimization of the acquisition function is limited to 100 seconds. This is due to the combination of the expensive MCMC procedure, and the large optimization formulation. We compare the fitting and optimization times against other methods in Fig. 8, and provide a further selection of wall-time comparisons in Section F.2.
We therefore recommend BARK for BO settings where the objective is expensive to evaluate, taking at least several minutes, or having a large associated financial cost. Here, the benefit of an accurate model with strong uncertainty quantification is worth the cost paid in wall-time. Note that the PestControl and MAX (material design) benchmarks reflect examples of such black-box functions. For settings where experiments are cheap and/or function evaluations are quick, we recommend alternate methods.
Figure 8: Time taken to fit the surrogate and optimize the acquisition function for the XGBoostMNIST benchmark.
@@ 8 Conclusion
Impact statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgments
We would like to thank Alexander Larionov for support in implementing benchmarks and baseline methods. Funding for this work was provided by BASF SE, Ludwigshafen am Rhein, and EPSRC through the Modern Statistics and Statistical Machine Learning CDT for TB (EP/Y034813/1) and JPF (EP/S023151/1). TB is also funded by the EPSRC IConIC Prosperity Partnership (EP/X025292/1). RM holds concurrent appointments as a Professor at Imperial and as an Amazon Scholar. This paper describes work performed at Imperial prior to joining Amazon and is not associated with Amazon.
Appendix A Related work comparison
We provide a summary of the existing literature in Table 2, that highlights key similarities and differences between the various BO methods covered in the literature review.
Table 2: Comparison of a selection of relevant existing approaches for Bayesian optimization. For each method, we describe the underlying surrogate used to model the objective, and the tool used to optimize the acquisition function. A ‘mixture’ kernel is the combination of a categorical kernel $k_{h}$ and a continuous kernel $k_{x}$ of the form $k(\mathbf{z},\mathbf{z^{\prime}})=k_{h}(\mathbf{h},\mathbf{h^{\prime}})\cdot k_{x}(\mathbf{x},\mathbf{x^{\prime}})+k_{h}(\mathbf{h},\mathbf{h^{\prime}})+k_{x}(\mathbf{x},\mathbf{x^{\prime}})$
Method Reference Surrogate Optimization GP Trees/ Partitions Mixed domain
BART Chipman et al. (2010) Bayesian sum-of-trees Grid
ENTMOOT Thebelt et al. (2021) Gradient boosted trees MIP
SMAC Lindauer et al. (2022) Random forest Random local search
LeafGP Thebelt et al. (2022) GP (Tree kernel) MIP
BARK Ours GP (Bayesian tree kernel) MIP
Exact GP Garrido-Merchán (2020) GP (Matérn) with input rounding Gradient descent
CASMO-POLITAN Wan et al. (2021) Local GP (Mixture) Alternate gradient descent + discrete perturbation
CoCaBo Ru et al. (2020) GP (Mixture) Multi-armed bandit + gradient descent
RKHS Embeddings Buathong et al. (2020) GP (Deep embeddings) Genetic optimization using derivatives
MiVaBo Daxberger et al. (2021) Linear model (with non-linear features) Thompson sampling
Gryffin Häse et al. (2021) BNN with simplex projection Gradient descent
HyBO Deshwal et al. (2021) GP (Diffusion kernel) CMA-ES: alternating continuous and discrete subspaces
PR Daulton et al. (2022) GP (Mixture) Gradient descent
TuRBO Eriksson et al. (2019) Local GP (Matérn) Thompson sampling
BAxUS Papenmeier et al. (2022) Local GP (Matérn, with embeddings) Thompson sampling
Tree structured Jenatton et al. (2017) GP (Matérn) + shared weights Path-based + gradient descent  [^2] This method requires prior knowledge of the structure of the tree function.
Appendix B Toy regression examples
As discussed in the main paper, our key motivation behind using tree-based kernels is the optimization of the acquisition function over mixed feature spaces. However, these models also model the underlying function differently, which we investigate in this section by considering some toy 1D problems in Fig. 9.
First, the tree kernel is better able to model non-stationary functions, where the lengthscale varies over the search space. For example, in Figs. 8(a) and 8(b) the very short lengthscale for small input values means that the RBF kernel learns a short lengthscale over the whole space, leading to large uncertainty around $x=0.5$, unlike the BARK model which has a much lower variance in this region.
Second, BARK has a prior belief about the correlation between categorical values. Since each split in the tree structure divides the categorical values into two subsets, the prior probability that two values will be placed in the same leaf is non-zero. Conversely, the indicator kernel used in Ru et al. (2020) explicitly assigns zero correlation between different categorical values. This results in increased variance for unseen categories, as shown in Fig. 8(c).
These properties are not necessarily ‘better’ - different black-box functions require different modeling assumptions. For example, the MAX benchmark in Section 7.3 requires fewer BO iterations when modeled using trees. However, other benchmarks are better modeled with the assumption of smoothness provided by Euclidean GP kernels.
(a) Continuous
(b) Discrete (ordered)
(c) Categorical (unordered)
Figure 9: Regression on a set of 1D toy functions, fit using a BARK GP and an RBF kernel. These functions highlight some notable differences in behavior between the two models.
Appendix C Notation
C.1 Table of symbols
@@
C.2 Model hyperparameters
For clarity, we provide an overview of the parameters used in the BARK model in Table 4.
Table 4: Hyperparameters and parameters of the BARK model. For fixed parameters, their value is given. Parameters that are sampled are given the label Sampled.
Symbol Description Value
$m$ Number of trees in ensemble 50
$(\alpha,\beta)$ Parameters of node depth prior $(0.95,2)$
$(\nu,q)$ Parameters of noise prior $(3,0.9)$
$\sigma_{0}^{2}$ Signal variance 1
$\sigma_{y}^{2}$ Noise variance Sampled
$T_{i}$ The $i$th tree structure Sampled
Appendix D Experimental details
D.1 Method details
BARK. For both model fitting and BO, we use 16 samples for the BARK kernel. Increasing the number of samples would (marginally) improve the model fit, however this would increase the complexity of the optimization problem, meaning that the optimization would take longer to converge. BARK uses 1000 burn-in samples and 400 kernel samples, running in parallel with 4 chains. We use a thinning rate of 100 to obtain the 16 samples. We evaluate the effect of increasing the number of samples in Appendix G.
@@ D.1 Method details
There is no procedure in the literature for optimizing acquisition functions for BART, and the MIP formulation would be too expensive for the 1000s of BART samples required. We therefore evaluate the surrogate on a grid of $\max(2^{5D},2^{14})$ points using a space-filling approach (Sobol, 1967). At each iteration, we sample a new grid to better explore the domain. This grid size was chosen to match the wall clock time of the BARK optimization (as shown in Fig. 11.
@@
We use the BoFire v0.0.16 implementation (Dürholt et al., 2024), which in turn uses BoTorch v0.11.3 (Balandat et al., 2020). To optimize mixed space acquisition functions, we use the optimize_acqf_mixed_alternating function provided by BoTorch.
@@
D.2 Benchmark details
A description of the benchmarks used in Bayesian optimization is given in Tables 5 and 6. For each BO benchmark, we generate a set of $\min(2D,30)$ initial points, and run the optimization loop for 100 iterations, reporting the minimum value observed up to iteration $t$.
@@ D.2 Benchmark details
Appendix E Additional results
E.1 Regression comparison with GP-RBF
In Section 7.1, we show that BART and BARK have similar modeling abilities, and that our kernel perspective on BART still leads to strong regression performance. For completeness, we also provide a comparison to the GP-RBF model, with linear combination of sum and product kernels, in Section E.1.
Table 7: Regression performance of the BART and GP-RBF models, given as mean (and standard deviation). The number of training points used for each benchmark is given by $n$. The number of dimensions for each problem is given by $D$ (continuous + integer + categorical). The better performing model is in bold for each metric.
Benchmark $n$ $D$ NLPD $\downarrow$ MSE $\downarrow$
BARK GP-RBF BARK GP-RBF
[…]
E.2 Continuous benchmarks
In the main experiments of the paper, we note the motivating use case for BARK is in optimizing over mixed feature spaces. However, we also provide two additional benchmarks here, for continuous synthetic problems. Specifically, we use the Hartmann and Styblinksi-Tang problems from the SFU library (Surjanovic & Bingham, 2013). We show the BO performance in Fig. 10. The smooth functions are more suited to GP-RBF, however BARK outperforms all other tree-based methods, and still shows good performance in this unfavorable setting on the Styblinski-Tang benchmark.
Figure 10: Bayesian optimization performance on benchmarks with a continuous-only domain: Hartmann (6D) and Styblinski-Tang (10D), both from the SFU library of synthetic functions.
Appendix F Time comparison
F.1 Posterior sampling
BART and BARK both take significantly longer to train than the other baselines, as they are MCMC-based methods. Given $N$ observations, training BART has $O(N)$ complexity, whereas BARK trains with $O(N^{3})$ complexity (which is reduced to $O(N^{2})$ when sampling tree structures due to the low-rank update discussed in Section 5). We therefore expect BART to scale far more favorably with many observations. In Section F.1, we show the time taken to train the models on the regression problems from Section 7. The fitting time for BARK is reasonable in the low-data regime of BO, matching that of BART when only 40 training points are used. Since we apply our model in the BO setting, where the evaluation of the black-box function may take hours, or possibly even days, we argue that this time taken is not a significant limitation of the model. We also note that we use fewer samples for BARK (400 samples per parallel chain) compared to BART (1000 samples per chain).